Calculate the standard free-energy change for the reaction at 25 ∘C.

Refer to the list of standard reduction potentials.


2Au3+(aq)+3Mg(s)↽−−⇀2Au(s)+3Mg2+(aq)


Δ𝐺∘= kJ










1
Expert's answer
2022-04-28T02:37:03-0400

The reduction half cell reactions and E0 values are:

Au3+(aq) + 3e →Au (s)   E0 = + 1.52 V 

Mg(s) - 2e  → Mg2+(aq) E0 = -2.36 V 


In the reaction, Au is being oxidized and so the overall cell potential is:

E0 = ((-2.36) – (+1.52)) V = -3.88 V


The reaction involves 6 electrons so, using ΔG0 = -nFE0

ΔG0 = - (6) × (96485 C mol-1) × (-3.88 V) =2246171 J mol-1 = 2.25 × 103 kJ mol-1